### v1

# A Graph Neural Network Approach for Mapping the Conceptual Structure and Inter-Branch Connectivity of Physics

Abstract
The vast and interconnected body of physical law represents a complex network of knowledge whose higher-order structure is not always explicit. This work introduces a novel framework for representing and analyzing physical laws as a comprehensive, weighted knowledge graph. We constructed a database of 659 distinct physical equations, which was subjected to a rigorous semantic cleaning process to resolve notational ambiguities, resulting in a high-fidelity corpus of 400 advanced physics equations, intentionally excluding elementary mechanics to focus on inter-branch connections in modern physics. We developed an enhanced graph representation where both physical concepts and equations are nodes, connected by weighted inter-equation bridges. These weights are objectively defined using normalized metrics for variable overlap, physics-informed importance scores, and bibliometric data. A Graph Attention Network (GAT), with hyperparameters optimized via grid search, was trained for link prediction. The model achieved a test AUC of $0.9742\pm 0.0018$ across five independent 5000 epoch runs (patience 500), significantly outperforming both classical heuristics (best baseline AUC: 0.9487, common neighbors) and established GNN architectures like GraphSAGE (AUC: $0.9504\pm 0.0128$, $p=0.029$). Statistical testing confirmed the significance of all comparisons ($p<0.05$), with the model showing a 2.7% improvement over the best baseline. Our analysis reveals three key findings: (i) The model autonomously rediscovers the known macroscopic structure of physics, identifying strong conceptual axes between Electromagnetism and Statistical Mechanics. (ii) It identifies central “hub” equations that serve as critical bridges between multiple physical domains. (iii) The model generates stable, computationally-derived hypotheses for cross-domain relationships, identifying both known principles and suggesting novel mathematical analogies for further theoretical investigation. The framework can generate hundreds of such hypotheses, enabling the creation of specialized datasets for targeted analysis of specific physics subfields.
1 Introduction
The accumulated knowledge of physics comprises a vast corpus of mathematical equations. These laws are traditionally organized into distinct branches—such as Classical Mechanics, Electromagnetism, and Quantum Mechanics. While this categorization is useful, it can obscure the deeper, structural similarities that form a “syntactic grammar” underlying physical theory. Identifying these hidden connections is crucial, as historical breakthroughs have often arisen from recognizing analogies between seemingly disparate fields [1, 2].
A significant challenge in the computational analysis of scientific knowledge is notational polysemy, where a single symbol can represent different concepts (e.g., $P$ for Pressure or Power). This ambiguity can create spurious connections and confound statistical analyses [3, 4]. In recent years, Graph Neural Networks (GNNs) have emerged as a powerful tool for analyzing complex relational data [5, 6]. Graph Attention Networks (GATs) [7, 8], in particular, are well-suited for knowledge graph analysis [9, 10].
Recent advances in machine learning for scientific discovery have demonstrated the potential for automated hypothesis generation [11, 12]. Link prediction methods have proven particularly effective for knowledge graph completion [13, 14], making them ideal for discovering latent mathematical analogies in scientific databases.
I hypothesize that the entire body of physical law can be modeled as a network, where a rigorously validated GNN can identify statistically significant structural patterns. This paper presents a methodology to build and analyze such a graph. My objectives are threefold:
  1. 1.
    To develop a robust pipeline for converting a symbolic database of physical laws into a semantically clean, enhanced, weighted knowledge graph, with edge weights derived from objective, literature-based metrics.
  2. 2.
    To train and statistically validate a parsimonious GNN to learn the structural relationships within this graph and predict plausible, non-trivial connections.
  3. 3.
    To analyze the model’s top-ranked predictions to map the syntactic connectivity of physics and generate novel, testable hypotheses about shared mathematical formalisms.
This preliminary work suggests that this approach can serve as a powerful engine for scientific inquiry.
2 Methods
2.1 Dataset Curation and Semantic Disambiguation
The foundation of this work is a curated database of 659 physical laws, compiled from academic sources into a JSON format and parsed using SymPy [17]. A crucial preliminary step was semantic disambiguation to resolve notational polysemy, a well-documented challenge in physics equation databases [15, 16]. My systematic disambiguation process identified and resolved 213 ambiguous equations out of 639 total equations (33.3% modification rate). The algorithmic scan flagged variables appearing across conceptually distant physics domains using quantitative criteria: variables were disambiguated when they appeared in $\geq 3$ distinct physics branches with different semantic meanings. The context-sensitive renaming protocol applied domain-specific suffixes (e.g., q $\rightarrow$ q_charge in electromagnetism vs. q_heat in thermodynamics) and standardized fundamental constants (e.g., _light $\rightarrow$ c, h_planc_lightk_boltzmann $\rightarrow$ h). Table 1 summarizes the most frequent variable corrections and their cross-domain distribution.
Table 1: Top Variable Disambiguations and Cross-Domain Analysis
Variable Correction Frequency Affected Domains
qr $\rightarrow$ sqrt 40 QM, Modern Physics, Classical Mechanics
ome_chargega $\rightarrow$ omega 29 Electromagnetism, Statistical Mechanics
lamda $\rightarrow$ lambda 20 Optics, Quantum Mechanics
_light $\rightarrow$ c 19 11 domains (most frequent constant)
q $\rightarrow$ q_charge/q_heat 16 Electromagnetism, Thermodynamics
P $\rightarrow$ P_power 10 Electromagnetism, Thermodynamics
Cross-Domain Variable Statistics
Total shared variables across domains 109
Variables in $\geq 5$ domains 25 High disambiguation priority
Most ubiquitous: c, t, m 80, 74, 70 Universal physics constants
After semantic cleaning, I obtained a corpus of 657 equations. Elementary mechanics laws were intentionally excluded to focus the analysis on inter-branch connections in modern physics.
2.2 Enhanced Knowledge Graph Construction
The cleaned dataset was transformed into a weighted, undirected graph where nodes represent either equations or physical concepts. The edge weight formula, a critical component of my methodology, was designed to be objective and mathematically sound, incorporating three normalized components:
$w_{ij}=\alpha\cdot J(V_{i},V_{j})+\beta\cdot I(V_{i},V_{j})+\gamma\cdot S(B_{i},B_{j})$ (1)
where:
  • $J(V_{i},V_{j})$ is the Jaccard Index for variable overlap, $\frac{|V_{i}\cap V_{j}|}{|V_{i}\cup V_{j}|}$, ensuring a normalized measure of syntactic similarity.
  • $I(V_{i},V_{j})$ represents a normalized, physics-informed importance score for the shared variables, where fundamental constants (e.g., $c,h,k_{B}$) are assigned higher a priori weights based on their centrality in physical theory.
  • $S(B_{i},B_{j})$ is the continuous branch similarity, derived from quantitative scientometric studies that map the structure of science based on citation patterns, such as the work of Börner et al. [23].
The hyperparameters $\alpha,\beta,\gamma$ were constrained to sum to 1 and the configuration was identified through grid and sensitivity analyses spanning the simplex of admissible weights, systematically evaluating dozens of candidates across all statistical metrics of interest. Results demonstrated that weight combinations in the vicinity of the values selected consistently yielded best performance, minimal variance, and stable behavior under cross-validation and perturbation, confirming the methodological robustness of the fusion scheme and the absence of overfitting to any specific parameter setting. All tested configurations were benchmarked with ablation and alternative combination methods as recommended by current best practices [24, 25]. The optimal values were found to be $\alpha=0.5$, $\beta=0.35$, and $\gamma=0.15$. This configuration although it yields a slightly lower AUC compared to the previous optimum, could promote the discovery of a broader and more diverse set of cross-domain connections. Increasing the weight of syntactic similarity ($\alpha$) reduces bias toward established pathways encoded by the importance and branch similarity terms, thereby facilitating the identification of less obvious mathematical analogies that might be overlooked by more “conservative” parameter settings.
2.3 Model Architecture and Training
To address concerns of potential overfitting raised by the peer review, I designed a parsimonious Graph Attention Network (“Super GNN”) with a significantly reduced parameter count.
Figure 1: Super GNN training and prediction analysis. (Top left) Training loss evolution showing stable convergence. (Top right) Validation metrics (AUC and Average Precision) over 0.96. (Bottom left) Precision at different recall levels demonstrating excellent performance across all operating points. (Bottom right) Score distribution histogram with 90th percentile at 0.97, indicating highly confident predictions.
  • Architecture: A 3-layer GAT utilizing GATv2Conv layers [8] with hidden dimensions of 64 $\rightarrow$ 32 $\rightarrow$ 16.
  • Attention Heads: A decreasing multi-head attention mechanism of 4 $\rightarrow$ 2 $\rightarrow$ 1.
  • Total Parameters: The final model has only 52,065 trainable parameters, ensuring a reasonable parameter-to-data ratio for graph size. The training dynamics and model performance are illustrated in Figure 1.
3 Results and Analysis
To better assess the model’s discriminative power, I performed a statistical validation on the full sample of predicted links (see Figure 2). Artificial negative controls (n=7,200n=7{,}200) were generated, and the FDR was estimated via bootstrap (500 iterations) at multiple thresholds. For the recommended threshold (score $\geq 0.90$), the mean FDR was $0.001$ [95% CI: 0.0000.0000.0050.005] with 3,1003{,}100 predictions above this cutoff. However, rare negative controls with scores $\geq 0.9$ (n=0n=022 per run) highlight residual false positive risk. The analysis identifies a critical intermediate zone (score 0.60.60.850.85, 7,8777{,}877 predictions) where the estimated noise proportion reaches up to 15%. These results support robust but cautious use of hard thresholds and call for further expert validation of top predictions.
Figure 2: Distribution of prediction scores (n=79,800) and artificial negative controls (n=7,200), FDR curve with bootstrap-derived 95% confidence intervals, and sensitivity heatmap for varying noise/signal definitions. High confidence predictions (score $\geq 0.90$, n=3,100) show FDR $\approx 0.1\%$ [0–0.5%], but up to 2 negative controls exceed this threshold in some runs. The intermediate zone (score 0.6–0.85, n=7,877) shows a potential noise overlap of 10–15%. These findings demonstrate overall strong separation, but also caution that rare high-score false positives and ambiguous intermediate cases require careful downstream validation.
3.0.1 Neural Architecture Comparisons with Statistical Testing
Statistical significance testing across 5 independent runs revealed significant performance differences: SUPER GNN vs. GraphSAGE ($0.9742\pm 0.0018$ vs. $0.9504\pm 0.0128$, $p=2.90\times 10^{-2}$), vs. GCN (0.9742 vs. $0.9364\pm 0.0090$, $p=1.70\times 10^{-3}$), and vs. simplified GAT (0.9742 vs. $0.9324\pm 0.0161$, $p=5.78\times 10^{-3}$). The SUPER GNN achieved a 2.7% improvement over the best classical baseline (common neighbors: 0.9487). The ablation study demonstrated that removing edge weights reduced performance to AUC = 0.9306, while using single attention heads yielded AUC = 0.8973, validating the importance of the architectural design choices including the physics-aware decoder with bilinear component, edge weight utilization, and multi-head attention mechanism.
Table 2: Model Performance Comparison with Statistical Validation
Method Mean AUC $\pm$ Std vs. SUPER GNN p-value
Classical Baselines
Common Neighbors 0.9487 -2.55% $9.00\times 10^{-6}$
Adamic-Adar 0.9481 -2.61% -
Jaccard Index 0.9453 -2.89% -
Preferential Attachment 0.8728 -10.14% -
Neural Methods (Identical $64\rightarrow 32\rightarrow 16$ Architecture)
SUPER GNN (Full) $0.9742\pm 0.0018$ - -
GraphSAGE $0.9504\pm 0.0128$ -2.38% $2.90\times 10^{-2}$
GCN $0.9364\pm 0.0090$ -3.78% $1.70\times 10^{-3}$
GAT (Simplified) $0.9324\pm 0.0161$ -4.18% $5.78\times 10^{-3}$
Ablation Studies
No Edge Weights 0.9306 -4.36% -
Single Attention Head 0.8973 -7.69% -
3.1 The Rediscovered Macrostructure of Physics
A “Cross-Branch Connection Strength” heatmap confirms the model’s ability to reproduce the known structure of physics, identifying strong conceptual axes like Thermodynamics $\leftrightarrow$ Statistical Mechanics and Electromagnetism $\leftrightarrow$ Optics. The ego network analysis revealed striking patterns in my knowledge graph of 644 nodes (400 equations, 244 concepts) connected by 12,018 edges. Figure 3 provides a comprehensive overview of the most connected fundamental concepts, showing their individual network topologies and revealing the hierarchical structure of physics knowledge.
Figure 3: Overview of fundamental physics concepts showing ego networks for 12 key variables. Each subplot displays the network topology around concepts like Energy (E, 12 equations), Force (F, 18 equations), speed of light (c, 52 equations), and Planck constant (h, 29 equations). The density and structure of connections reveal the centrality and cross-domain importance of each concept, with variables like ’c’ and ’m’ showing the densest interconnection patterns.
Figure 4 demonstrates my methodological approach to identifying the most significant connections by filtering for equation-equation bridges with strong connections, resulting in a high-quality network of 224 connections among 132 equations.
As shown in Figure 5, the variable ’q’ emerged as particularly interesting, revealing notational ambiguity across 26 equations split between Electromagnetism (charge) and Thermodynamics (heat), demonstrating the framework’s capability for automated data quality assessment.
The variable ’k’ emerged as the most connected concept with 43 equations across 6 branches, confirming its role as a fundamental constant appearing in diverse contexts—from Coulomb’s constant in electromagnetism to Boltzmann’s constant in thermodynamics. The visualization shows a dense, almost circular pattern of connections, reflecting its mathematical ubiquity.
Figure 4: Network visualization of strong equation-equation bridges, showing 224 connections among 132 equations with a density of 0.0259. This filtered network reveals the strongest mathematical relationships while maintaining clear clustering by physics branch. The varying edge thickness represents connection strength, and the professional layout demonstrates my rigorous filtering approach for identifying meaningful connections.
Figure 5: Ego network for variable ’q’ revealing notational ambiguity with 26 connections split between Electromagnetism (21 equations, charge) and Thermodynamics (5 equations, heat). This visualization led to the discovery of polysemy in dataset, where ’q’ represents both electric charge and heat flow, demonstrating the framework’s capability for automated data quality assessment.
3.2 Computationally-Generated Cross-Domain Hypotheses
The graph neural network framework identified several high-scoring links between physical equations from different domains. These connections represent computationally-generated hypotheses about a shared mathematical syntax underlying physics. From the stable connections that persisted across multiple random seeds and experimental runs, we selected eight that appeared particularly intriguing from a theoretical physics perspective, either for their conceptual novelty or for their validation of known principles through purely data-driven methods. The framework can generate hundreds of such hypotheses, enabling the creation of specialized cross-domain datasets for targeted theoretical investigations and deeper mathematical analysis of specific physics subfields. The following selection highlights the framework’s dual capability: rediscovering fundamental principles of modern physics and identifying novel, non-trivial mathematical analogies. Among the 143 stable connections identified, Statistical Mechanics emerges with 93 connections as the central unifying branch, while approximately 70% of connections occur between different physics domains.
3.2.1 Debye Length and Dirac Equation (Score: 0.9691)
The model identifies an intriguing connection between the Debye screening length in plasma physics (adv_0146: $\lambda_{D}=\sqrt{\epsilon_{0}kT/(ne^{2})}$) and the Dirac equation for relativistic fermions (adv_0179: $(i\gamma^{\mu}\partial_{\mu}-mc/\hbar)\psi=0$). This links classical collective phenomena with relativistic quantum mechanics. The Debye length describes how charges in a plasma collectively screen electric fields over a characteristic distance, while the Dirac equation governs the behavior of relativistic electrons. This connection suggests potential theoretical insights for relativistic plasma physics in extreme astrophysical environments.
3.2.2 Hydrostatic Pressure and Maxwell Distribution (Score: 0.9690)
A conceptually significant connection was found between the hydrostatic pressure formula (adv_0116: $P=P_{0}+\rho gh$) and the Maxwell-Boltzmann velocity distribution (adv_0160: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$). This links a macroscopic formula with the statistical distribution of molecular velocities, revealing how macroscopic pressure emerges from the microscopic velocity distribution of molecular collisions.
3.2.3 Maxwell Distribution and Adiabatic Process (Score: 0.9687)
A significant connection was identified between the Maxwell-Boltzmann (adv_0160) distribution: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$) and the adiabatic process for ideal gases (bal_0129: $PV^{\gamma}=\text{constant}$). This connection reveals the relationship between the statistical distribution of molecular velocities and the thermodynamic behavior of gases under adiabatic conditions. The framework identifies how microscopic velocity distributions directly determine macroscopic thermodynamic properties during rapid compressions or expansions.
3.2.4 Doppler Effect and Four-Momentum (Score: 0.9685)
The framework identified an analogy between the Doppler effect for sound waves (adv_0130: $f^{\prime}=f_{0}\sqrt{\frac{v+v_{r}}{v-v_{s}}}$) and the relativistic four-momentum invariant (adv_0176: $p_{\mu}p^{\mu}=(mc)^{2}$). This connection identifies mathematical similarities between frequency transformations in media and energy-momentum transformations in spacetime, suggesting possibilities for acoustic analogues of relativistic phenomena in exotic media.
3.2.5 Blackbody Radiation and Terminal Velocity (Score: 0.9680)
An unconventional connection links Planck’s blackbody radiation (adv_0094: $B(\lambda,T)=\frac{8\pi hc}{\lambda^{5}(e^{hc/\lambda kT}-1)}$) with terminal velocity from fluid dynamics (bal_0239: $v_{t}=\sqrt{2mg/(\rho AC_{d})}$). The model identifies that radiation pressure from stellar emission can balance gravitational forces, creating an astrophysical terminal velocity where photon pressure acts analogously to fluid resistance—demonstrating the framework’s ability to identify non-obvious cross-domain mathematical structures.
3.2.6 Radioactive Decay and Induced EMF (Score: 0.9651)
The model identified a structural analogy between radioactive decay (adv_0061: $N(t)=N_{0}e^{-\lambda t}$) and electromagnetic induction (bal_0027: $\mathcal{E}=-L\frac{dI}{dt}$). Both phenomena are governed by first-order differential equations with exponential solutions. This mathematical isomorphism between nuclear physics and electromagnetism exemplifies the framework’s ability to uncover purely syntactic analogies independent of physical mechanisms.
3.2.7 Blackbody Radiation and Navier-Stokes (Score: 0.9611)
A connection was found between blackbody radiation (adv_0095: $B(f,T)=\frac{2hf^{3}}{c^{2}(e^{hf/kT}-1)}$) and the Navier-Stokes equation (bal_0245). This suggests the possibility of treating photon gases as fluids with pressure and viscosity properties, identifying radiation hydrodynamics as a potential research direction for understanding cosmic background radiation evolution and photon-matter interactions.
3.2.8 Spherical Capacitance and Fine Structure Constant (Score: 0.9624)
The model identified a subtle connection between spherical capacitance (adv_0014: $C=4\pi\epsilon_{0}R$) and the fine structure constant (bal_0219: $\alpha=e^{2}/(4\pi\epsilon_{0}\hbar c)$). Both formulations contain the fundamental term $4\pi\epsilon_{0}$, revealing how the permittivity constant governs both macroscopic electrostatics and quantum electromagnetic interaction strength, demonstrating a fundamental unity between classical and quantum electromagnetism.
These findings demonstrate the framework’s capability to identify both established physical principles and novel mathematical analogies. The consistent identification of connections between Statistical Mechanics and other domains validates the framework’s ability to uncover the mathematical structures that unify disparate areas of physics.
Figure 6 shows the relationship between neural predictions and embedding similarity, revealing a moderate correlation (0.394) that suggests the model learns complex, non-linear relationships beyond simple vector similarity.
Figure 6: Neural prediction scores vs embedding cosine similarity for 400 equation pairs. The moderate correlation (0.394) indicates the GNN learns complex relationships beyond simple vector similarity. High-scoring predictions (top right) represent the most confident cross-domain discoveries, while the distribution across similarity values shows the model’s ability to identify connections even between mathematically dissimilar equations.
3.3 Validation Through Negative Results and Methodological Controls
Equally informative as the model’s successful predictions are its “failures” and the spurious connections it learned to avoid. These negative results provide crucial validation, demonstrating that the framework is sensitive to genuine mathematical formalism rather than arbitrary correlations or philosophical links. We highlight three key types of such results.
  • Elimination of Spurious Links via Disambiguation: A critical validation of our methodology is demonstrated by the model’s behavior before and after semantic disambiguation. When trained on the raw, non-disambiguated graph, the model produced numerous high-confidence but physically nonsensical predictions. A prime example was a top-ranked link (Score >0.98>0.98) between AC Power (Electromagnetism) and the Adiabatic Process (Thermodynamics). The model correctly identified a strong mathematical link based on the shared variable ‘P’, but it was blind to the fact that ‘P’ represented Power in the first equation and Pressure in the second.
    After our semantic disambiguation protocol was applied—systematically renaming variables like ‘P’ to ‘P_power’ and ‘P_pressure’—this high-scoring false positive was entirely eliminated from the results. This demonstrates that our disambiguation step is not merely a preprocessing formality but a crucial component of the framework. It acts as an effective “auditor” of scientific notation, ensuring the final discoveries are grounded in genuine mathematical similarity rather than notational coincidence.
  • Absence of Purely Conceptual Links: The model correctly avoids predicting links between concepts that are deeply related physically but syntactically dissimilar. For instance, the connection between Maxwell’s equations and Special Relativity—a cornerstone of modern physics—received a low score. This is expected, as their standard formulations share almost no symbolic variables. This “failure” correctly demonstrates that the framework operates as a syntactic pattern recognizer, not a system with latent physical knowledge.
  • Low Scores for Distant Formalisms: Finally, the model correctly assigns low scores to pairs of equations that are conceptually linked but mathematically distinct. The connection between Newton’s Laws and the Schrödinger Equation, for example, scored low (0.621). This reflects not only their different mathematical structures but also aligns with our explicit exclusion of elementary mechanics from the final dataset to focus the model on the formalisms of modern physics. Similarly, the profound analogy between Thermodynamic and Information Entropy received only a moderate score (0.743), indicating the model cannot capture this deep connection without an explicit information-theoretic context.
These results, both the elimination of false positives and the presence of true negatives, validate that the model’s discoveries are grounded in the mathematical syntax of the provided equations, not in arbitrary correlations or a misunderstanding of the underlying physics.
4 Discussion
The results may support my hypothesis that a graph neural network, when combined with a methodologically sound graph representation, can identify statistically significant mathematical patterns in a corpus of physics equations. The model’s ability to achieve a robust and high AUC score ($0.9742\pm 0.0030$) after a semantic disambiguation and validation process indicates the discovered relationships may reflect true structural similarities in the mathematical formalism, not notational artifacts or model overfitting.
The discovered connections must be interpreted as syntactic analogies, not as direct physical or causal relationships. For instance, the link between Photon Energy (Quantum Mechanics) and the relativistic Four-Momentum invariant (Relativity) does not imply a new physical theory connecting them. Rather, it suggests a “meta-law”: a shared mathematical grammar used to describe critical phenomena, even when they occur in vastly different energy regimes.
The model is “semantically blind”— it recognizes syntax, not physical causality. Therefore, its predictions must be interpreted as hypotheses that require further theoretical investigation.
5 Conclusion and Future Work
I have developed and preliminarily validated a GNN-based framework capable of mapping the mathematical structure of physical law. Our proposed GAT model achieves state-of-the-art performance (AUC = 0.9742) on this novel task, demonstrably outperforming classical and established GNN baselines in a statistically-validated comparison. By addressing key methodological challenges—from semantic disambiguation to objective edge weight definition—we have shown that this approach can identify non-trivial, physically interesting, and statistically significant patterns.
5.1 Current Limitations and Future Directions
  • Database: The immediate next step is to expand the equation database to over 1,000 equations. This richer corpus will not only broaden the map but also enable the encoding of deeper mathematical structures—such as the presence of specific operators, derivatives, or symmetries—as features, moving the analysis from a syntactic to a more profound structural level.
  • Dynamic Methodological Refinements: Future work will explore more adaptive methodologies. This includes developing learnable edge-weighting schemes that go beyond a fixed linear combination and extending our disambiguation protocol to act as a general-purpose “auditor” for standardizing notational conventions across different scientific knowledge bases.
  • Advanced Statistical Analysis: Building on the rigorous FDR validation presented, future research can focus on more nuanced metrics that assess the novelty and non-triviality of discoveries. We also plan to explore methods from causal inference to investigate the potential directionality of the identified mathematical relationships.
  • Interactive Tools for Expert Validation: To bridge the gap between computational patterns and physical insight, we plan to develop interactive visualization tools that facilitate an expert-in-the-loop process to validate and interpret the most promising machine-generated analogies.
5.2 Broader Implications
This work opens several avenues for the broader scientific community, including applications in education, philosophy of science, and other scientific domains like chemistry and biology.
6 Data and Code Availability
All code, the cleaned dataset, and model weights will be available at: https://github.com/kingelanci/graphysics.
Correspondence
Correspondence should be addressed to Massimiliano Romiti (massimiliano.romiti@acm.org).
Acknowledgments
The author thanks the arXiv endorsers and the anonymous reviewers for their valuable feedback. Special gratitude to the ACM community for professional development resources.